OpenStudy (anonymous):

medal & fan!! simplify these plez 1+2(3-5) (-8)^2-2(6^2-9)

OpenStudy (texaschic101):

1 + 2(3 - 5) = 1 + 2(-2) = 1 - 4 = -3 ============= (-8)^2 - 2(6^2 - 9) = 64 - 2(36 - 9) = 64 - 2(27) = 64 - 54 = 10 ============= any questions ?

OpenStudy (anonymous):

2 more questions and im done... 6-(4^22-7)^2 8-12y^3/4

OpenStudy (texaschic101):

I am sorry....I am not good at these

OpenStudy (anonymous):

its okay lol